# Nonstandard Universes

### v1

@@ 3.3 ${}^{\sigma}A$ is External if $A$ is Infinite
In this subsection, we assume that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard framework for a set $X$, i.e. a transfer map with the property that there is a countable set $C\in\mathbb{U}$ such that ${}^{\sigma}C\subsetneqq{}^{*}C$.
@@
If $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard framework, then ${}^{\sigma}A$ is external whenever $A\in\mathbb{U}$ is infinite. Hence ${}^{\sigma}A\subsetneqq{}^{*}A$ whenever $A\in\mathbb{U}$ is infinite.
@@
This does not assert that ${}^{*}\!\!\preceq$ is a well-ordering on ${}^{*}C$, however, because it may not be the case that ${}^{*}\mathcal{P}(C)=\mathcal{P}({}^{*}C)$. What it does assert is that every non-empty internal subset of ${}^{*}C$ has a ${}^{*}\!\!\preceq$-least element, since we know that ${}^{*}\mathcal{P}(C)=\mathcal{P}({}^{*}C)\cap{}^{*}\mathbb{U}$ is the set of internal subsets of ${}^{*}C$. In particular, if $D$ is internal, there is a ${}^{*}\!\!\preceq$-least element $d_{0}$ of $D$.
Now for every $n\in\mathbb{N}$,
@@
$\mathbb{V}\vDash\forall x\in{}^{*}C\;(x={}^{*}c_{0}\lor x={}^{*}c_{1}\lor\dots\lor x={}^{*}c_{n}\lor x\>{}^{*}\!\!\succ{}^{*}c_{n})$
@@
Now since ${}^{*}C$ is internal, it follows that ${}^{\sigma}C$ is external, because the difference of two internal sets is internal.
Now suppose that $A\in\mathbb{U}$ is infinite, and that $f\in\mathbb{U}$ is a surjection $f:A\twoheadrightarrow C$. It follows easily that ${}^{*}f:{}^{*}A\twoheadrightarrow{}^{*}C$. Since ${}^{*}(f(a))={}^{*}f({}^{*}a)$ for all $a\in A$, we have ${}^{*}f[{}^{\sigma}A]={}^{\sigma}C$. If ${}^{\sigma}A$ is internal, then by Lemma 3.6 it would follows that ${}^{\sigma}C$ is internal as well. As this is false, ${}^{\sigma}A$ is external.
@@ 4.1 The Set ${}^{*}\mathbb{N}$ of Hypernatural Numbers
Throughout this section, assume that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard embedding for an infinite set $X$. Without loss of generality (e.g. by renaming elements) we may assume that $\mathbb{N}\subseteq X$. Then as $\mathcal{P}(X)\in\mathbb{U}$ and $\mathbb{U}$ is transitive, also $\mathbb{N}\in\mathbb{U}$. Since ${}^{*}x=x$ for all $x\in X$, it follows in particular that ${}^{\sigma}\mathbb{N}=\mathbb{N}$. Since * is nonstandard, $\mathbb{N}\subsetneqq{}^{*}\mathbb{N}$.
@@
shows that any member of ${}^{*}\mathbb{N}-\mathbb{N}$ is $>N$. As this is true for all $N\in\mathbb{N}$, it follows that every member of ${}^{*}\mathbb{N}-\mathbb{N}$ is greater than any natural number, and thus said to be infinite. We thus define $\mathbb{N}_{\infty}:={}^{*}\mathbb{N}-\mathbb{N}$ to be the set of infinite natural numbers, so that we have the disjoint union ${}^{*}\mathbb{N}=\mathbb{N}\cup\mathbb{N}_{\infty}$. As $\mathbb{N}={}^{\sigma}\mathbb{N}$ is external, so is $\mathbb{N}_{\infty}$.
Now the structure $(\mathbb{N},+,\dotdiv,\cdot,0,1)\in\mathbb{U}$ is elementarily equivalent to the corresponding structure $({}^{*}\mathbb{N},+,\dotdiv,\cdot,0,1)\in{}^{*}\mathbb{U}$ (where $n\dotdiv m:=n-m$ if $n\geq m$ and $:=0$ else). We should really have written $({}^{*}\mathbb{N},{}^{*}+,{}^{*}\dotdiv,{}^{*}\cdot,0,1)$, but we drop the stars on the arithmetic operations for easier reading. It is clear that if $n\in\mathbb{N}_{\infty}$ then $n-1\in\mathbb{N}_{\infty}$ also, and thus $\mathbb{N}_{\infty}$ has no least element, i.e. ${}^{*}\mathbb{N}$ is not well-ordered. However:
@@
    (Overflow) Let $N\in\mathbb{N}$, and suppose that $X\in{}^{*}\mathcal{P}(\mathbb{N})$ is an internal subset of ${}^{*}\mathbb{N}$ with the property that whenever $n\in\mathbb{N}$ satisfies $n\geq N$, then $n\in X$. Then there is $M\in\mathbb{N}_{\infty}$ such that whenever $n\in{}^{*}\mathbb{N}$ satisfies $N\leq n\leq M$, then $n\in X$
@@
    (Underflow) Let $M\in\mathbb{N}_{\infty}$, and suppose that $X\in{}^{*}\mathcal{P}(\mathbb{N})$ is an internal subset of ${}^{*}\mathbb{N}$ with the property that whenever $n\in{}^{*}\mathbb{N}_{\infty}$ satisfies $n\leq M$, then $n\in X$. Then there is $N\in\mathbb{N}$ such that whenever $n\in{}^{*}\mathbb{N}$ satisfies $N\leq n\leq M$, then $n\in X$
Proof: (a) By the internal definition principle, the set $Y:=\{n\in{}^{*}\mathbb{N}:n>N\land n\in{}^{*}\mathbb{N}-X\}$ is internal. If $Y=\varnothing$, then every $M\in\mathbb{N}_{\infty}$ has the desired property. If $Y$ is non-empty, then it must have a least element $K$, i.e. $K$ is least such that $K>N$ and $K\not\in X$. By the assumption on $X$ we must have $K\in\mathbb{N}_{\infty}$. It follows that every infinite natural number which is $<K$ belongs to $X$, so let $M:=K-1$.
@@ 4.2 Hyperfinite sets
In this section, we suppose that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map over a set $X$, where $\mathbb{N}\subseteq X$. Recall that the $<$-relation on $\mathbb{N}$ is just a subset of $<\;\subseteq\mathbb{N}\times\mathbb{N}$, so to say that $m<n$ is equivalent to saying $(m,n)\in\;<$.
@@
In that case, we denote #$\#A=n$.
@@
Observe that if $f:\{0,1,\dots,n-1\}\to A$ is an internal bijection, then $A=\text{\rm ran}(f)$ is internal, i.e. every hyperfinite set is internal.
@@
Then $\Psi(A,f,n)$ is a bounded formula which asserts that $f$ is a bijection $f:\{0,\dots,n-1\}\to A$. Transfer guarantees that ${}^{*}\Psi(A,f,n)$ asserts the same for $A,f\in{}^{*}\mathbb{U}$ and $n\in{}^{*}\mathbb{N}$. We thus see that, by definition, $A\in\mathbb{V}$ is hyperfinite if and only if there are an internal $f$ and an $n\in{}^{*}\mathbb{N}$ such that ${}^{*}\Psi(A,f,n)$. Suppose that $B\in\mathbb{U}$, and recall that ${}^{*}\mathcal{P}(B)=\mathcal{P}({}^{*}B)\cap{}^{*}\mathbb{U}$ is the set of all internal subsets of $B$. Let
@@
Proof: ()($\Rightarrow$)( ⇒ ): Suppose that $A\in\mathbb{V}$ is hyperfinite, and let $f:\{0,1,\dots,n-1\}\to A$ be an internal bijection, for some $n\in{}^{*}\mathbb{N}$. Let $U:=\{0,1,\dots,n-1\}=\text{\rm dom}(f)$ and $A:=\text{\rm ran}(f)$. Then $U,A$ are internal so $U\in{}^{*}\mathcal{P}(\mathbb{N})$, and there is a transitive set $B\in\mathbb{U}$ such that $A\in{}^{*}B$. As ${}^{*}B$ is transitive also, we have $A\subseteq{}^{*}B$, and since $f$ is internal, we must have $f\in\mathcal{P}({}^{*}\mathbb{N}\times{}^{*}B)\cap{}^{*}\mathbb{U}={}^{*}\mathcal{P}(\mathbb{N}\times B)$. It follows that ${}^{*}\Psi(A,f,n)$ holds. Now observe that
@@
()($\Leftarrow$)( ⇐ ): Conversely, suppose that $A\in{}^{*}\mathcal{P}^{<\omega}(B)$ for some $B\in\mathbb{U}$. Observe that
@@ 5.1 Definitions and Basic Properties
Throughout this section, assume that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$.
@@
    $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement if and only if for every set $\Sigma(x)$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{\mathbb{U}}$, if $\Sigma(x)$ is finitely satisfiable in $\mathbb{U}$ by elements of some set $T\in\mathbb{U}$, then $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$.
@@
    $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated if and only if for every set $\Sigma(x)$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$, if $\Sigma(x)$ is finitely satisfiable in $\mathbb{V}$ by elements of some set $T\in{}^{*}\mathbb{U}$, then $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$.
@@
    $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement if and only if whenever $\mathcal{A}\subseteq\mathbb{U}$ is a family of sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p., then $\bigcap{}^{\sigma}\mathcal{A}=\bigcap\{{}^{*}A:A\in\mathcal{A}\}$ is non-empty.
@@
    $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated if and only if whenever $\mathcal{A}\subseteq{}^{*}\mathbb{U}$ is a family of internal sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p., then $\bigcap\mathcal{A}$ is non-empty.
@@
Proof: (a) Suppose first that $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement, and that $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}\subseteq\mathbb{U}$ is a family of sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p. By replacing each $A_{\beta}$ by $A_{\beta}\cap A_{0}$ we may without loss of generality assume that each $A_{\beta}\subseteq A_{0}$ — This does not affect the f.i.p. nor the intersection of all the $A_{\beta}$. Let $\Sigma(x):=\{\sigma_{\beta}(x):\beta<\alpha\}$, where $\sigma_{\beta}(x)\equiv x\in A_{\beta}$. Then $\Sigma(x)$ is a set of $<\kappa$-many bounded formulas of $\mathcal{L}_{\mathbb{U}}$ which is finitely satisfiable by elements of $A_{0}$. Hence $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$. Thus there is $a\in{}^{*}\mathbb{U}$ such that $a\in{}^{*}A_{\beta}$ for all $\beta<\alpha$, so that $\bigcap{}^{\sigma}\mathcal{A}\neq\varnothing$.
@@
For the reverse direction, suppose we have a set $\Sigma(x)=\{\sigma_{\beta}(x):\beta<\alpha\}$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{\mathbb{U}}$, and that $\Sigma(x)$ is finitely satisfiable in $\mathbb{U}$ by elements of some set $T\in\mathbb{U}$. Let $A_{\beta}:=\{t\in T:\mathbb{U}\vDash\sigma_{\beta}(t)\}$. As $\Sigma(x)$ is finitely satisfiable, the family $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}$ has the f.i.p. If $s\in\bigcap{}^{\sigma}\mathcal{A}$, then $s\in{}^{*}\mathbb{U}$. As $\mathbb{U}\vDash\forall x\in A_{\beta}(\sigma_{\beta}(x))$, we have $\mathbb{V}\vDash\forall x\in{}^{*}A_{\beta}(\sigma_{\beta}(x))$ so that $\mathbb{V}\vDash\Sigma(s)$.
@@
(b) Suppose first that $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated, and that $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}\subseteq{}^{*}\mathbb{U}$ is a family of internal sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p. By replacing each $A_{\beta}$ by $A_{\beta}\cap A_{0}$ we may without loss of generality assume that each $A_{\beta}\subseteq A_{0}$ — This does not affect the f.i.p. nor the intersection of all the $A_{\beta}$. Let $\Sigma(x):=\{\sigma_{\beta}(x):\beta<\alpha\}$, where $\sigma_{\beta}(x)\equiv x\in A_{\beta}$. Then $\Sigma(x)$ is a set of $<\kappa$-many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$ which is finitely satisfiable by elements of $A_{0}$. Hence $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$. Thus there is $a\in{}^{*}\mathbb{U}$ such that $a\in A_{\beta}$ for all $\beta<\alpha$, so that $\bigcap\mathcal{A}\neq\varnothing$.
@@
For the reverse direction, suppose we have a set $\Sigma(x)=\{\sigma_{\beta}(x):\beta<\alpha\}$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$, and that $\Sigma(x)$ is finitely satisfiable in $\mathbb{V}$ by elements of some set $T\in{}^{*}\mathbb{U}$. Let $A_{\beta}:=\{t\in T:\sigma_{\beta}(t)\}$. By the internal definition principle, each $A_{\beta}$ is internal. As $\Sigma(x)$ is finitely satisfiable by members of $T$, the family $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}$ has the f.i.p. If $s\in\bigcap\mathcal{A}$, then $s\in{}^{*}\mathbb{U}$, and $\mathbb{V}\vDash\Sigma(s)$.

### v2

@@ 3.3 $^{\sigma}A$ is External if $A$ is Infinite
In this subsection, we assume that $\mathbb{U}\to\mathbb{V}$ is a nonstandard framework for a set $X$, i.e. a transfer map with the property that there is a countable set $C\in\mathbb{U}$ such that $^{\sigma}C\subsetneqq{}^{*}C$.
@@
If $\mathbb{U}\to\mathbb{V}$ is a nonstandard framework, then $^{\sigma}A$ is external whenever $A\in\mathbb{U}$ is infinite. Hence $^{\sigma}A\subsetneqq{}^{*}A$ whenever $A\in\mathbb{U}$ is infinite.
@@
This does not assert that $^{*}\!\!\preceq$ is a well-ordering on $^{*}C$, however, because it may not be the case that $^{*}\mathcal{P}(C)=\mathcal{P}({}^{*}C)$. What it does assert is that every non-empty internal subset of $^{*}C$ has a ${}^{*}\!\!\preceq$-least element, since we know that $^{*}\mathcal{P}(C)=\mathcal{P}({}^{*}C)\cap{}^{*}\mathbb{U}$ is the set of internal subsets of $^{*}C$.
Now suppose that $D$ is internal, and let $d_{0}$ be the ${}^{*}\!\!\preceq$-least element of $D$. We shall obtain a contradiction.
Observe that, for every $n\in\mathbb{N}$,
@@
$\mathbb{V}\vDash\forall x\in{}^{*}C\;(x={}^{*}c_{0}\lor x={}^{*}c_{1}\lor\dots\lor x={}^{*}c_{n}\lor x\>{}^{*}\!\!\succ{}^{*}c_{n}).$
@@
But as $^{*}C$ is internal, it follows that $^{\sigma}C$ is external, because the difference of two internal sets is internal.
Suppose now that $A\in\mathbb{U}$ is infinite, and that $f\in\mathbb{U}$ is a surjection $f:A\twoheadrightarrow C$. It follows easily that $^{*}f:{}^{*}A\twoheadrightarrow{}^{*}C$. Since $^{*}(f(a))={}^{*}f({}^{*}a)$ for all $a\in A$, we have $^{*}f[{}^{\sigma}A]={}^{\sigma}C$. If $^{\sigma}A$ is internal, then by Lemma 3.6 it would follows that $^{\sigma}C$ is internal as well. As this is false, $^{\sigma}A$ is external.
@@ 4.1 The Set $^{*}\mathbb{N}$ of Hypernatural Numbers
Throughout this section, assume that $\mathbb{U}\to\mathbb{V}$ is a nonstandard embedding for an infinite set $X$. Without loss of generality (e.g. by renaming elements) we may assume that $\mathbb{N}\subseteq X$. Then as $\mathcal{P}(X)\in\mathbb{U}$ and $\mathbb{U}$ is transitive, also $\mathbb{N}\in\mathbb{U}$. Since $^{*}x=x$ for all $x\in X$, it follows in particular that $^{\sigma}\mathbb{N}=\mathbb{N}$. Since * is nonstandard, $\mathbb{N}\subsetneqq{}^{*}\mathbb{N}$.
@@
shows that any member of $^{*}\mathbb{N}-\mathbb{N}$ is >N>N. As this is true for all $N\in\mathbb{N}$, it follows that every member of $^{*}\mathbb{N}-\mathbb{N}$ is greater than any natural number, and thus said to be infinite. We thus define $\mathbb{N}_{\infty}:={}^{*}\mathbb{N}$-$\mathbb{N}$ to be the set of infinite natural numbers, so that we have the disjoint union $^{*}\mathbb{N}=\mathbb{N}\cup\mathbb{N}_{\infty}$. As $\mathbb{N}={}^{\sigma}\mathbb{N}$ is external, so is $\mathbb{N}_{\infty}$.
Now the structure $(\mathbb{N},+,\dotdiv,\cdot,\leq,0,1)\in\mathbb{U}$ is elementarily equivalent to the corresponding structure $({}^{*}\mathbb{N},+,\dotdiv,\cdot,\leq,0,1)\in{}^{*}\mathbb{U}$ (where $n\dotdiv m:=n-m$ if $n\geq m$ and :=0:=0 else). We should really have written $({}^{*}\mathbb{N},{}^{*}+,{}^{*}\dotdiv,{}^{*}\cdot,{}^{*}\leq,0,1)$, but we drop the stars on the arithmetic operations and order relation for easier reading. It is clear that if $n\in\mathbb{N}_{\infty}$ then $n-1\in\mathbb{N}_{\infty}$ also, and thus $\mathbb{N}_{\infty}$ has no least element, i.e. $^{*}\mathbb{N}$ is not well-ordered. However:
@@
    (Overflow) Let $N\in\mathbb{N}$, and suppose that $X\in{}^{*}\mathcal{P}(\mathbb{N})$ is an internal subset of $^{*}\mathbb{N}$ with the property that whenever $n\in\mathbb{N}$ satisfies $n\geq N$, then $n\in X$. Then there is $M\in\mathbb{N}_{\infty}$ such that whenever $n\in{}^{*}\mathbb{N}$ satisfies $N\leq n\leq M$, then $n\in X$ .
@@
    (Underflow) Let $M\in\mathbb{N}_{\infty}$, and suppose that $X\in{}^{*}\mathcal{P}(\mathbb{N})$ is an internal subset of $^{*}\mathbb{N}$ with the property that whenever $n\in{}^{*}\mathbb{N}_{\infty}$ satisfies $n\leq M$, then $n\in X$. Then there is $N\in\mathbb{N}$ such that whenever $n\in{}^{*}\mathbb{N}$ satisfies $N\leq n\leq M$, then $n\in X$ .
Proof: (a) By the internal definition principle, the set $Y:=\{n\in{}^{*}\mathbb{N}:n>N\land n\in{}^{*}\mathbb{N}-X\}$ is internal. If $Y=\varnothing$, then every $M\in\mathbb{N}_{\infty}$ has the desired property. If $Y$ is non-empty, then it must have a least element $K$, i.e. $K$ is least such that $K>N$ and $K\not\in X$. By the assumption on $X$ we must have $K\in\mathbb{N}_{\infty}$. It follows that every infinite natural number which is <K<K belongs to $X$, so let M:=K1M:=K-1.
@@ 4.2 Hyperfinite sets
In this section, we suppose that $\mathbb{U}\to\mathbb{V}$ is a transfer map over a set $X$, where $\mathbb{N}\subseteq X$. Recall that the <<-relation on $\mathbb{N}$ is just a subset of $\subseteq\mathbb{N}\times\mathbb{N}$, so to say that $m<n$ is equivalent to saying $(m,n)\in\;<$.
@@
In that case, we denote #A=n\#A=n.
@@
Observe that if $n\in{}^{*}\mathbb{N}$, then $\{0,1,\dots,n-1\}=\{x\in{}^{*}\mathbb{N}:x<n\}$ is internal, by the internal definition principle. Hence if $f:\{0,1,\dots,n-1\}\to A$ is an internal bijection, then $A=\text{\rm ran}(f)$ is internal, i.e. every hyperfinite set is internal.
@@
Then $\Psi(A,f,n)$ is a bounded formula which asserts that $f$ is a bijection $f:\{0,\dots,n-1\}\to A$. Transfer guarantees that $^{*}\Psi(A,f,n)$ asserts the same for $A,f\in{}^{*}\mathbb{U}$ and $n\in{}^{*}\mathbb{N}$. We thus see that, by definition, $A\in\mathbb{V}$ is hyperfinite if and only if there are an internal $f$ and an $n\in{}^{*}\mathbb{N}$ such that $^{*}\Psi(A,f,n)$.
Suppose that $B\in\mathbb{U}$, and recall that $^{*}\mathcal{P}(B)=\mathcal{P}({}^{*}B)\cap{}^{*}\mathbb{U}$ is the set of all internal subsets of $B$. Let
@@
Proof: $(\Rightarrow)$: Suppose that $A\in\mathbb{V}$ is hyperfinite, and let $f:\{0,1,\dots,n-1\}\to A$ be an internal bijection, for some $n\in{}^{*}\mathbb{N}$. Let $U:=\{0,1,\dots,n-1\}=\text{\rm dom}(f)$ and $A:=\text{\rm ran}(f)$. Then $U,A$ are internal so $U\in{}^{*}\mathcal{P}(\mathbb{N})$, and there is a transitive set $B\in\mathbb{U}$ such that $A\in{}^{*}B$. As $^{*}B$ is transitive also, we have $A\subseteq{}^{*}B$, and since $f$ is internal, we must have $f\in\mathcal{P}({}^{*}\mathbb{N}\times{}^{*}B)\cap{}^{*}\mathbb{U}={}^{*}\mathcal{P}(\mathbb{N}\times B)$. It follows that $^{*}\Psi(A,f,n)$ holds. Now observe that
@@
$(\Leftarrow)$: Conversely, suppose that $A\in{}^{*}\mathcal{P}^{<\omega}(B)$ for some $B\in\mathbb{U}$. Observe that
@@ 5.1 Definitions and Basic Properties
Throughout this section, assume that $\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$.
@@
    $\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement if and only if for every set $\Sigma(x)$ of $\kappa-$many bounded formulas of $\mathcal{L}_{\mathbb{U}}$, if $\Sigma(x)$ is finitely satisfiable in $\mathbb{U}$ by elements of some set $T\in\mathbb{U}$, then $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of $^{*}\mathbb{U}$.
@@
    $\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated if and only if for every set $\Sigma(x)$ of $\kappa-$many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$, if $\Sigma(x)$ is finitely satisfiable in $\mathbb{V}$ by elements of some set $T\in{}^{*}\mathbb{U}$, then $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of $^{*}\mathbb{U}$.
@@
    $\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement if and only if whenever $\mathcal{A}\subseteq\mathbb{U}$ is a family of sets of cardinality $\kappa$ such that $\mathcal{A}$ has the f.i.p., then $\bigcap{}^{\sigma}\mathcal{A}=\bigcap\{{}^{*}A:A\in\mathcal{A}\}$ is non-empty.
@@
    $\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated if and only if whenever $\mathcal{A}\subseteq{}^{*}\mathbb{U}$ is a family of internal sets of cardinality $\kappa$ such that $\mathcal{A}$ has the f.i.p., then $\bigcap\mathcal{A}$ is non-empty.
@@
Proof: (a) Suppose first that $\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement, and that $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}\subseteq\mathbb{U}$ is a family of sets of cardinality $\kappa$ such that $\mathcal{A}$ has the f.i.p. By replacing each $A_{\beta}$ by $A_{\beta}\cap A_{0}$ we may without loss of generality assume that each $A_{\beta}\subseteq A_{0}$ — This does not affect the f.i.p. nor the intersection of all the $A_{\beta}$. Let $\Sigma(x):=\{\sigma_{\beta}(x):\beta<\alpha\}$, where $\sigma_{\beta}(x)\equiv x\in A_{\beta}$. Then $\Sigma(x)$ is a set of $\kappa-$many bounded formulas of $\mathcal{L}_{\mathbb{U}}$ which is finitely satisfiable by elements of $A_{0}$. Hence $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of $^{*}\mathbb{U}$. Thus there is $a\in{}^{*}\mathbb{U}$ such that $a\in{}^{*}A_{\beta}$ for all $\beta<\alpha$, so that $\bigcap{}^{\sigma}\mathcal{A}\neq\varnothing$.
@@
For the reverse direction, suppose we have a set $\Sigma(x)=\{\sigma_{\beta}(x):\beta<\alpha\}$ of $\kappa-$many bounded formulas of $\mathcal{L}_{\mathbb{U}}$, and that $\Sigma(x)$ is finitely satisfiable in $\mathbb{U}$ by elements of some set $T\in\mathbb{U}$. Let $A_{\beta}:=\{t\in T:\mathbb{U}\vDash\sigma_{\beta}(t)\}$. As $\Sigma(x)$ is finitely satisfiable, the family $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}$ has the f.i.p. If $s\in\bigcap{}^{\sigma}\mathcal{A}$, then $s\in{}^{*}\mathbb{U}$. As $\mathbb{U}\vDash\forall x\in A_{\beta}(\sigma_{\beta}(x))$, we have $\mathbb{V}\vDash\forall x\in{}^{*}A_{\beta}(\sigma_{\beta}(x))$ so that $\mathbb{V}\vDash\Sigma(s)$.
@@
(b) Suppose first that $\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated, and that $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}\subseteq{}^{*}\mathbb{U}$ is a family of internal sets of cardinality $\kappa$ such that $\mathcal{A}$ has the f.i.p. By replacing each $A_{\beta}$ by $A_{\beta}\cap A_{0}$ we may without loss of generality assume that each $A_{\beta}\subseteq A_{0}$ — This does not affect the f.i.p. nor the intersection of all the $A_{\beta}$. Let $\Sigma(x):=\{\sigma_{\beta}(x):\beta<\alpha\}$, where $\sigma_{\beta}(x)\equiv x\in A_{\beta}$. Then $\Sigma(x)$ is a set of $\kappa-$many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$ which is finitely satisfiable by elements of $A_{0}$. Hence $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of $^{*}\mathbb{U}$. Thus there is $a\in{}^{*}\mathbb{U}$ such that $a\in A_{\beta}$ for all $\beta<\alpha$, so that $\bigcap\mathcal{A}\neq\varnothing$.
@@
For the reverse direction, suppose we have a set $\Sigma(x)=\{\sigma_{\beta}(x):\beta<\alpha\}$ of $\kappa-$many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$, and that $\Sigma(x)$ is finitely satisfiable in $\mathbb{V}$ by elements of some set $T\in{}^{*}\mathbb{U}$. Let $A_{\beta}:=\{t\in T:\sigma_{\beta}(t)\}$. By the internal definition principle, each $A_{\beta}$ is internal. As $\Sigma(x)$ is finitely satisfiable by members of $T$, the family $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}$ has the f.i.p. If $s\in\bigcap\mathcal{A}$, then $s\in{}^{*}\mathbb{U}$, and $\mathbb{V}\vDash\Sigma(s)$.